OpenStudy (anonymous):

How to convert sec^2(x)sin(2x) to 2tan(x)

OpenStudy (anonymous):

\[\sec ^2 (x)\sin(2x) = 2\tan(x) \]

OpenStudy (anonymous):

I know that \[1+\tan^2(x) = \sec^2(x)\]

OpenStudy (anonymous):

So i have \[(1+\tan^2(x))(\sin(2x))\]

OpenStudy (anonymous):

then sin(2x) = 2sin(x)cos(x) so I have \[(1+\tan^2(x))(2\sin(x)\cos(x))\]

OpenStudy (unklerhaukus):

use this double angle formula \[\sin(2\theta)=2\sin(\theta)\cos(\theta)\] like you have already, but instead of doing that to \(\sec^2\), use \(\sec=1/\cos\)

OpenStudy (anonymous):

Ahhhhhhh ok so I get 1/cos^2

OpenStudy (unklerhaukus):

yeah, and you can now cancel a cos

OpenStudy (anonymous):

... my damn brain... I thought that that would not work because it was to simple

OpenStudy (anonymous):

Thanks man.

OpenStudy (unklerhaukus):

thats alright, these thing are tricky, i usually try to convert every term to sin(x) and cos(x) first

OpenStudy (anonymous):

sounds good, now how would I do \[\tan(x)2\cos(2x)\]

OpenStudy (anonymous):

\[\frac{ \sin(x) }{ \cos(x) }2(one of the three...)\]

OpenStudy (unklerhaukus):

\(one with a \cos (x) in it\)

OpenStudy (anonymous):

cool man, got that now... goodness gracious me I hope this type of differentiation equation does not pop up in my exam... the orrigional equation was to find the first derivative of \[y = \frac{ 1 }{ 2 }\tan(x)\sin(2x)\]

OpenStudy (unklerhaukus):

the original equation?

OpenStudy (unklerhaukus):

just apply the double angle formula and convert the tan to sines and cosines